OpenStudy (help108):

The work of a student to solve the equation 3(3y − 8) = 10 + 3y is shown below: Step 1: 3(3y − 8) = 10 + 3y Step 2: 6y − 5 = 10 + 3y Step 3: 3y − 5 = 10 Step 4: 3y = 15 Step 5: y = 5 In which step did the student first make an error and what is the correct step?

OpenStudy (photon336):

multiply the 3 by everything in the ( ) first

OpenStudy (help108):

so step 1 is wrong?

OpenStudy (anonymous):

yes

OpenStudy (photon336):

3(3y-8) = 3*3y-3*8

OpenStudy (photon336):

that's what we mean by using the distributive property

OpenStudy (photon336):

This would give us the following: 9y-24 = 10+3y now the idea here would be to get the y's on the same side.

OpenStudy (help108):

so which step is wrong?

OpenStudy (help108):

@Photon336

OpenStudy (photon336):

Step one is wrong

OpenStudy (anthonyym):



OpenStudy (help108):

thanks

OpenStudy (help108):

which step is correct? @Photon336

OpenStudy (help108):

hello? @Photon336

OpenStudy (help108):

@Photon336

OpenStudy (help108):

@Photon336

OpenStudy (help108):

@Photon336

OpenStudy (anthonyym):

If you mean if 3*3y - 3*8 is the same as (3*3y) - 3*8, yes it is. Using order of operations, you would still simplify to 9y - 24

OpenStudy (photon336):

Because the first step is wrong the others are too.

OpenStudy (photon336):

we need to start with 9y-24 = 10+3y

OpenStudy (help108):

The work of a student to solve the equation 3(3y − 8) = 10 + 3y is shown below: Step 1: 3(3y − 8) = 10 + 3y Step 2: 6y − 5 = 10 + 3y Step 3: 3y − 5 = 10 Step 4: 3y = 15 Step 5: y = 5 In which step did the student first make an error and what is the correct step? Step 2: 9y − 24 = 10 + 3y Step 2: 6y + 11 = 10 + 4y Step 3: 6y − 8 = 3(6 + y) Step 3: 3(3y − 8) = 7 + y

OpenStudy (help108):

i put options now is it B or A?

OpenStudy (help108):

hurry dude i just need this one awnser i want to sleep plz @Photon336

OpenStudy (help108):

@Photon336

OpenStudy (help108):

@Photon336

OpenStudy (help108):

@Photon336

OpenStudy (anonymous):

step 1 is wrong

OpenStudy (help108):

i know but which one is right..............

OpenStudy (anonymous):

ok let me think

OpenStudy (anonymous):

idk man srry

OpenStudy (help108):

:(* ok

OpenStudy (anthonyym):

So do you know what step 2 should be?